# Fine-Grained Safety Neurons with Training-Free Continual Projection to Reduce LLM Fine Tuning Risks

### v3

@@ Abstract
Fine-tuning as service injects domain-specific knowledge into large language models (LLMs), while challenging the original alignment mechanisms and introducing safety risks. A series of defense strategies have been proposed for the alignment, fine-tuning, and post-fine-tuning phases, where most post-fine-tuning defenses rely on coarse-grained safety layer mapping. These methods lack a comprehensive consideration of both safety layers and fine-grained neurons, limiting their ability to efficiently balance safety and utility. To address this, we propose the Fine-Grained Safety Neurons (FGSN) with Training-Free Continual Projection method to reduce the fine-tuning safety risks. FGSN inherently integrates the multi-scale interactions between safety layers and neurons, localizing sparser and more precise fine-grained safety neurons while minimizing interference with downstream task neurons. We then project the safety neuron parameters onto safety directions, improving model safety while aligning more closely with human preferences. Extensive experiments across multiple fine-tuned LLM models demonstrate that our method significantly reduce harmfulness scores and attack success rates with minimal parameter modifications, while preserving the model’s utility. Furthermore, by introducing a task-specific, multi-dimensional heterogeneous safety neuron cluster optimization mechanism, we achieve continual defense and generalization capability against unforeseen emerging safety concerns.
Introduction
Existing large language models (LLMs) have demonstrated remarkable capabilities and have been widely applied in various aspects of production and daily life, including language understanding (y Arcas 2022), mathematical reasoning (Shao et al. 2024), and domains such as healthcare (Liu et al. 2025b) and finance (Hu et al. 2025). However, the safety concerns associated with LLMs are also expanding, encompassing privacy (Kibriya et al. 2024), child safety (Kurian 2024), animal protection (Kanepajs et al. 2025), pharmaceuticalcon regulations (Hakim et al. 2024), political sensitivity (Goodman 2024), and violence (Liu et al. 2025a), and have drawn significant attention from regulators and enterprises across various countries (Fang and Perkins 2024; Hakim et al. 2024). Notably, even LLMs that have undergone safety alignment through supervised fine-tuning (SFT) (Devlin et al. 2018) and reinforcement learning from human feedback (RLHF) (Stiennon et al. 2020) may still experience disruptions during downstream task fine-tuning. This risk can be triggered by as few as less than 1% intentionally or unintentionally mixed harmful fine-tuning samples, or even by fine-tuning on entirely benign data, which may cause the model to forget its original safety alignment (Qi et al. 2023). Concurrently, safety concerns continue to evolve, making it a necessity for LLMs to maintain both downstream task utility and support continual safety improvements for broader adoption.
@@ Introduction
Existing safety defenses for fine-tuned downstream tasks can be categorized into three stages: preference alignment, downstream task fine-tuning, and post-fine-tuning safeguard mechanisms. RLHF preference alignment defenses (Huang, Hu, and Liu 2024; Rosati et al. 2024; Tamirisa et al. 2024)add perturbations during training to improve robustness to harmful queries and reduce safety forgetting during fine-tuning, but their lack of specificity causes performance instability across safety scenarios. Fine-tuning-stage (Bianchi et al. 2023; Zong et al. 2024; Huang et al. 2024b)incorporate safety data into the fine-tuning dataset for joint training, resulting in additional training overhead. Post-fine-tuning defenses  (Hsu et al. 2024; Casper et al. 2024; Huang et al. 2024a) separate LLM training/fine-tuning from safety protection and do not require backpropagation, achieving safety enhancement by editing part of the model’s weights or hidden states. Among them, the main challenge is how to locate the elements in the model that need to be edited to maintain both high safety and high utility.
@@
Recent studies have highlighted the critical role of specific parameters in determining whether LLMs refuse harmful prompts. For example,  Li et al. (2024) localize safety-critical layers, while SafeLoRA (Hsu et al. 2024) and Jailbreak Antidote (Shen et al. 2024) modify weights and hidden states within each layer, respectively. NLSR (Yi et al. 2025) identifies safety directions tied to matrix rank. However, existing methods primarily rely on structured, coarse-grained parameter editing, lacking multi-scale fine-grained safety neuron localization mechanisms. Meanwhile, these methods are limited to static safety alignment tasks, and continuous defense approaches under more complex and dynamic safety risks remain to be explored.
@@
In this paper, we propose Fine-Grained Safety Neurons with Training-Free Continual Projection to address safety risks caused by LLM’s fine-tuning. The proposed method introduces adaptive integration localization of multi-scale safety layers and neurons, which more precisely selects fewer safety-related parameters for safety projection, achieving enhanced safety performance while eliminating interference with generalization. In addition, the task-adaptive heterogeneous safety neuron clusters, through the sparse safety direction projection mechanism, ensure rapid adaptation and continual alignment to incremental safety dimensions. Our primary contributions are as follows:
  • We propose a multi-scale fine-grained method for safety neuron identification and training-free continual alignment, which adapts to localize safety neurons based on inter-layer safety importance analysis, minimizing interference with general-purpose neurons. Training-free projection of sparse safety neurons ensures efficient defense while preserving LLM utility.
@@
Related Work
@@
Safety Fine-Tuning of LLMs
@@
Existing external black-box defense methods suppress harmful outputs through complex filters or prompt engineering. For instance, xxx employs perplexity-based detection to reduce harmful outputs (Alon and Kamfonas 2023). The Self-Defense (Phute et al. 2023) method uses the model itself as a discriminator, converting harmful content into a refusal mode when detected. Self-reminders (Xie et al. 2023) enhance model safety awareness by adding system prompts (Xie et al. 2023), while the ICA (Wei et al. 2023) incorporates safety refusal examples before the prompt (Wei et al. 2023). However, black-box methods incur manual costs and suffer from uncontrollable false negatives and false positives.
@@
Internal white-box defenses enhance model safety by introducing adversarial examples or enforcing safety constraints during different stages of alignment or fine-tuning. Vaccine (Huang, Hu, and Liu 2024) adds random perturbations to each layer during alignment, but its performance is limited due to the lack of targeted safety considerations. SafeInstr (Bianchi et al. 2023) integrates safety alignment data during the fine-tuning process, while Goal Priority (Zhang et al. 2023) combines a refusal discriminator with fine-tuning to reclassify harmful outputs. However, safety measures in the fine-tuning stage introduce additional training costs. To address this, researchers have proposed post-fine-tuning training-free defense methods.
Post-Fine-Tuning Safety Defense
@@
Training-free post-fine-tuning defense techniques employ heuristic methods to add hidden state biases or modify weights. For instance, Jailbreak Antidote (Shen et al. 2024) applies safety bias corrections to activations based on transformer layers, while SafetyLock (Zhu et al. 2024) adjusts the output bias of attention heads.  Li et al. (2024) identifies safety-critical layers by analyzing activation differences between aligned and unaligned LLMs. SafeLoRA (Li et al. 2024) modifies the weights of layers with low similarity by comparing model weights before and after safety projection. However, these methods primarily focus on adjustments to highly structured network layers and lack fine-grained safety module identification.
@@
Wei et al. (2024) evaluates neuron-level safety relevance, pruning neurons that are highly responsive to unsafe prompts, while Antidote (Huang et al. 2024a) prunes harmful weights. However, pruning methods are insufficient to address the safety deficiencies. And these approaches focus only on local safety importance, lacking a comprehensive consideration of the internal relationships between safety layers and safety neurons, which limits their ability to balance persistent safety and generalizability.
Method
@@
Addressing safety concerns in fine-tuned models, we aim to achieve high defense success rates with minimal parameter modifications under a training-free setting, building the capability for continuous safety improvement. To achieve this, we design a multi-scale fine-grained safety neuron localization method and perform safety-direction projection on the sparse safety neuron weights, as shown in Fig. 1.
Figure 1: The fine-grained safety neurons with training-free continual projection framework. Our method consists of three main stages: downstream task fine-tuning, fine-grained safety neuron identification, and training-free continual projection.
Safety-Critical Layer Identification
@@
Considering that different layers in LLMs have varying impacts on safety alignment, we first theoretically analyzed the internal state differences across Transformer layers between the unaligned and aligned models, focusing on their responses to both benign and harmful prompts. Specifically, we presented the models with 100 benign prompts from the Alpaca instruction-following dataset (Taori et al. 2023) and 100 harmful prompts from the JailbreakBench safety dataset (Mazeika et al. 2024). Then, we systematically recorded the hidden states at each layer for both categories of prompts in five repeated experiments and compared the layer-wise outputs between a representative, extensively trained official base model and its instruction-aligned counterpart.
@@
Specifically, based on the LLaMA 3.1-8B (Meta AI 2024) model, we collected the mean hidden states at each layer under four conditions: 1) base model (LLaMA 3.1-8B) with benign prompts; 2) base model with harmful prompts; 3) instruction-aligned model (LLaMA 3.1-8B-instruct) with benign prompts; and 4) instruction-aligned model with harmful prompts. The aligned model is formally specified in Eq. 1, with the base model defined analogously.
@@
$\begin{array}[]{l}S_{\text{align}}(\text{benign})=\frac{1}{B}\sum_{b=1}^{B}D(o_{b}^{0},o_{b}^{1},\dots,o_{b}^{K-1})\\[6.0pt] S_{\text{align}}(\text{harm})=\frac{1}{H}\sum_{h=1}^{H}D(o_{h}^{0},o_{h}^{1},\dots,o_{h}^{K-1})\end{array}$ (1)
@@
To further analyze the impact ratio of safety alignment across the layers, we computed the cosine similarity between cases 1) and 2) to represent the base model, and between cases 3) and 4) to represent the aligned model, as shown in Eq. 2. Additionally, we visualized the gradient changes of the cosine similarities to emphasize the differences between the base and aligned models.
@@
As shown in Fig. 2a, both models exhibit a decreasing trend in cosine similarity across different types of prompts, indicating that the ability to distinguish harmful prompts increases progressively from shallow to deeper layers. Notably, significant differences emerge between the aligned model and the base model in layers 10 to 25, where the base model consistently exhibits higher cosine similarity values. This suggests that alignment enhances the model’s ability to differentiate harmful inputs in the mid-to-deep layers. Gradient analysis of cosine similarity (Fig. 2b) reveals that the most pronounced divergence occurs between layers 10 and 15, indicating that the internal representations of the two models begin to significantly diverge at approximately one-third of the model’s depth, leading to differing downstream behaviors.
Figure 2: Visualization of safety layers and safety neurons. (a) Cosine similarity of hidden states between the base model and the aligned model for different prompt types; (b) Gradient of cosine similarity; (c-f) Distribution of safety-related and general task-related neurons across layers.
Based on this observation, we identify layers 10–15 as safety-critical layers in the LLaMA model, with similar patterns observed in other models (see Appendix B for details). Within these safety-critical layers, we increase the threshold for fine-grained safety-relevant neuron identification to enable the projection of a broader set of safety neurons (as detailed in the next subsection).
Fine-Grained Safety Neuron Localization
After identifying the safety-critical layers, we incorporate multi-scale layer information into the fine-grained safety neuron localization, making it more sparse and precise. Specifically, we reintroduce the sampled benign and harmful datasets from the previous section, where harmful prompts elicit either safe responses. Fig. 2c-f visualizes the neurons with high activations (top 50%) for harmful and benign samples. Around one-third into the model depth, the two begin to gradually diverge, leading to significantly different responses in the deeper layers, further highlighting the interplay between multi-scale layers and neurons. Additionally, there is an overlap between the two, particularly in the shallow and mid-layers, necessitating the exclusion of general-purpose neurons while localizing safety neurons.
To comprehensively evaluate the importance of individual neurons within the model, we incorporate a data-driven neuron importance score based on activation states. Since our method is training-free and does not rely on backpropagation, the weights remain fixed. We use the sum of its input weights, serving as a proxy for its contribution. Formally, the safety-relevant neuron importance score $C_{j}^{harm}$ is computed as Eq. 3, and the importance score for general-task neurons $C_{j}^{benign}$ is computed in the same manner.
@@
$\text{Mask}_{l}[j]=\mathbb{I}\left(j\in\text{Top}_{q_{l}}(C^{\text{harm}})\setminus\text{Top}_{p_{l}}(C^{\text{benign}})\right)$ (4)
@@
Training-Free Sparse Projection
@@
Next, we project the identified sparse safety neurons onto the safety direction. This work employs the LoRA method (Hu et al. 2022) for efficient fine-tuning on downstream tasks. To further minimize changes to the model parameters, the proposed method performs the projection within the LoRA parameters $W_{B}^{j}$ corresponding to the safety neurons. The computation is formulated as follow:
@@
Continual Safety Neuron Projection
@@
Where $\mathbb{I}()$ denotes the indicator function.
Sparse heterogeneous projections maintain consistency in the safety direction, enhancing the safety of new dimensions while preventing catastrophic forgetting of previously adjusted safety dimensions.
Experiments
Experimental Settings
Model Method Edit Parameters GPT-4o Judger Llama3.1-405B Judger Keywords ASR AlpacaEval Winrate
@@
Our FGSN 4.67% 1.37 1.36 14% 54.52%
Table 1: Comparison of Safety and Utility Across Different Defense Methods on Alpaca-Finetuned Models.
Method Judger Score/ASR Params Acc
Lora
Finetune
GPT-4o 3.95 0% 54.20%
@@
Datasets and Models. To evaluate the effectiveness, we adopt Llama3.1-8B-Instruct (Meta AI 2024) and Qwen2.5-7B-Instruct (Yang et al. 2024) using the parameter-efficient fine-tuned LoRA approach on two distinct datasets: the instruction-following semantic Alpaca dataset (Taori et al. 2023) and the mathematical reasoning GSM8K dataset (Cobbe et al. 2021). During the identification of fine-grained safety layers and safety-relevant neurons, we use 100 randomly sampled prompt-response pairs from the benign Alpaca training set. The safety dataset comprises 100 harmful prompts from the “Goal” module of the JailbreakBench (Mazeika et al. 2024), along with safe responses generated by GPT-4o (OpenAI 2024). For defense evaluation, we assess overall safety performance using 100 randomly sampled general safety prompts from BeaverTails (Ji et al. 2023). This dataset also enables multi-dimensional safety evaluation to assess the continual alignment capabilities of FGSN. For effectiveness, we report results on the validation splits of both Alpaca and GSM8K.
Baseline Methods. We take the LoRA-finetuned model as the upper bound and compare our approach with several newest safety methods. Among them, self-reminders (Xie et al. 2023) are a prompt-based defense method through system-mode. Goal Priority (Zhang et al. 2023) combines discriminator enhancement and fine-tuning by transforming jailbreak prompts into safe refusals through targeted model tuning. SafeLoRA (Hsu et al. 2024) and Wanda (Wei et al. 2024) are both based on the same post-fine-tuning safety defense paradigm as ours, with SafeLoRA editing safety-relevant layers and Wanda targeting safety-relevant neurons.
Evaluation Metrics. We follow (Chao et al. 2024) to evaluate the proposed method from safety and utility. For safety, we adopt both LLM-based safety judger and keyword-based attack success rate (ASR). The LLM judger score ranges from 1 to 10, with higher scores indicating greater unsafe behavior. For utility, we use the AlpacaEval (Li et al. 2023) semantic evaluation benchmark to compute the WinRate between the original LoRA-finetuned model and the FGSN-modified model on the Alpaca dataset. We use accuracy to evaluate performance on the GSM8K dataset. Further experimental details are provided in Appendix A.
Performance on Static Safety Concerns
Enhancing safety defenses. In terms of safety judged by closed-source GPT-4o and open-source Llama3.1-405B, FGSN consistently achieves the lowest harmfulness scores across both Alpaca-finetuned models (1.02 / 1.27 on Llama-3-8B and 1.37 / 1.36 on Qwen-2.5-7B), clearly outperforming alternative methods such as Goal Priority ($\geq$2.10), Self Reminder ($\geq$2.01), and SafeLoRA ($\geq$1.51) as Tab. 1. Moreover, FGSN achieves the lowest keyword ASR on Alpaca-finetuned models (14% on both Llama and Qwen), demonstrating superior robustness compared to Goal Priority (35%, 30%) and Wanda (30%, 16%). These results indicate FGSN’s effectiveness in mitigating a broad spectrum of safety risks. Additionally, on the GSM8K-finetuned model, FGSN maintains strong performance, with GPT-4o and Llama3.1-405B harmfulness scores of 1.94 and 1.93, respectively—outperforming all other defense-capable baselines as Tab. 2.
Continual Safety Continual FGSN Judge Method Lora Finetune Universal Safety Animal Abuse Child Abuse Terrorism

### v4

# Multi-Level Safety Continual Projection for Fine-Tuned Large Language Models without Retraining

@@ Abstract
While fine-tuning services drive the rapid expansion of task capabilities in large language models (LLMs), they are often accompanied by the degradation and reorganization of safety-aligned representations, making models more prone to deviating from human preferences and exposing them to emerging jailbreak risks. Existing post-fine-tuning defense methods predominantly rely on single-scale safety correction mechanisms, which struggle to achieve a robust balance among safety, model utility, and continual adaptability. We propose Multi-Level Safety Continual Projection (MSCP), a training-free post-fine-tuning safety enhancement method that implicitly aligns global and localized safety activations through coordinated multi-level representations to isolate sparse neuron clusters governing safety-sensitive behaviors. It then applies composable safety-direction projections without retraining, effectively suppressing harmful outputs under minimal parameter perturbations while preserving task performance and improving alignment with human preferences. Extensive experiments across multiple fine-tuned LLM models demonstrate that our method significantly reduce harmfulness scores and attack success rates with minimal parameter modifications, while preserving the model’s utility. Furthermore, we introduce a task-specific, multi-dimensional heterogeneous safety activation clustering mechanism that enables continual defense and generalization capability against unforeseen emerging safety concerns.
1 Introduction
As large language models (LLMs) are continually adapted to specific tasks and application scenarios through fine-tuning Shao et al. (2024); Liu et al. (2025b); Hu et al. (2025), their safety boundaries are progressively reshaped during deployment Kibriya et al. (2024); Kurian (2024); Kanepajs et al. (2025); Liu et al. (2025a). Unlike static alignment assumptions, real-world safety requirements are inherently multi-dimensional, heterogeneous, and continually evolving, such that even models aligned via supervised fine-tuning (SFT) Devlin et al. (2018) and reinforcement learning from human feedback (RLHF) Stiennon et al. (2020) may experience degradation or forgetting of alignment representations during downstream task adaptation. Prior studies show that injecting only a small fraction of harmful samples—or even fine-tuning on entirely benign data—can trigger substantial post-fine-tuning safety degradationt Qi et al. (2023). This reality highlights the urgent need for post-fine-tuning safety mechanisms that preserve task utility while enabling training-free and continually adaptive safety enhancement.
@@ 1 Introduction
Existing safety defenses for fine-tuned downstream tasks can be categorized into three stages: preference alignment, downstream task fine-tuning, and post-fine-tuning safeguard mechanisms. RLHF preference alignment defenses Huang et al. (2024c); Rosati et al. (2024); Tamirisa et al. (2024)add perturbations during training to improve robustness to harmful queries and reduce safety forgetting during fine-tuning, but their lack of specificity causes performance instability across safety scenarios. Fine-tuning-stage Bianchi et al. (2023); Zong et al. (2024); Huang et al. (2024b)incorporate safety data into the fine-tuning dataset for joint training, resulting in additional training overhead. Post-fine-tuning defenses  Hsu et al. (2024); Casper et al. (2024); Huang et al. (2024a) separate LLM training/fine-tuning from safety protection and do not require backpropagation, achieving safety enhancement by editing part of the model’s weights or hidden states. Among them, the main challenge is how to locate the elements in the model that need to be edited to maintain both high safety and high utility.
@@
Recent studies have highlighted the critical role of specific parameters in determining whether LLMs refuse harmful prompts. For example,  Li et al. (2025) localize safety-critical layers, while SafeLoRA Hsu et al. (2024) and Jailbreak Antidote Shen et al. (2024) modify weights and hidden states within each layer, respectively. NLSR Yi et al. (2025) identifies safety directions tied to matrix rank. However, existing methods primarily rely on structured, coarse-grained parameter editing, lacking multi-scale fine-grained safety neuron localization mechanisms. Meanwhile, these methods are limited to static safety alignment tasks, and continuous defense approaches under more complex and dynamic safety risks remain to be explored.
@@
In this paper, we propose Multi-Level Safety Continual Projection (MSCP) to address safety risks caused by LLM’s fine-tuning. From a multi-scale representational perspective, MSCP captures the coordination between coarse-grained layers and fine-grained neurons in safety alignment, adaptively identifying sparse safety-critical parameter subsets and applying training-free safety-direction projections to enhance safety without compromising model utility. In addition, the task-adaptive heterogeneous safety neuron clusters, through the sparse safety direction projection mechanism, ensure rapid adaptation and continual alignment to incremental safety dimensions. Our primary contributions are as follows:
  • We propose a training-free continual safety alignment method based on coordinated multi-level representations, which implicitly aligns global and localized safety activations to identify sparse activation clusters governing safety-critical behaviors, and applies composable safety-direction projections, thereby achieving effective safety mitigation while preserving overall model utility.
@@
Figure 1: The multi-level safety continual projection for fine-tuned LLMs framework. Our method consists of three main stages: downstream task fine-tuning, multi-scale safety activation localization, and training-free continual projection.
2 Related Work
@@
2.1 Safety Fine-Tuning of LLMs
@@
Existing external black-box defense methods suppress harmful outputs through complex filters or prompt engineering. For instance, Alon and Kamfonas (2023) employs perplexity-based detection to reduce harmful outputs. The Self-Defense Phute et al. (2023) method uses the model itself as a discriminator, converting harmful content into a refusal mode when detected. Self-reminders Xie et al. (2023) enhance model safety awareness by adding system prompts Xie et al. (2023), while the ICA Wei et al. (2023) incorporates safety refusal examples before the prompt Wei et al. (2023). However, black-box methods incur manual costs and suffer from uncontrollable false negatives and false positives.
@@
Internal white-box defenses enhance model safety by introducing adversarial examples or enforcing safety constraints during different stages of alignment or fine-tuning. Vaccine Huang et al. (2024c) adds random perturbations to each layer during alignment, but its performance is limited due to the lack of targeted safety considerations. SafeInstr Bianchi et al. (2023) integrates safety alignment data during the fine-tuning process, while Goal Priority Zhang et al. (2023) combines a refusal discriminator with fine-tuning to reclassify harmful outputs. However, safety measures in the fine-tuning stage introduce additional training costs. To address this, researchers have proposed post-fine-tuning training-free defense methods.
2.2 Post-Fine-Tuning Safety Defense
@@
Training-free post-fine-tuning defense techniques employ heuristic methods to add hidden state biases or modify weights. For instance, Jailbreak Antidote Shen et al. (2024) applies safety bias corrections to activations based on transformer layers, while SafetyLock Zhu et al. (2024) adjusts the output bias of attention heads.  Li et al. (2025); Djuhera et al. (2025) identifies safety-critical layers by analyzing activation differences between aligned and unaligned LLMs. SafeLoRA Hsu et al. (2024) modifies the weights of layers with low similarity by comparing model weights before and after safety projection. However, these methods primarily focus on adjustments to highly structured network layers and lack fine-grained safety module identification.
@@
Wei et al. (2024) evaluates neuron-level safety relevance, pruning neurons that are highly responsive to unsafe prompts, while Antidote Huang et al. (2024a); Ao et al. (2025); Ao and Rumchurn (2025) prunes harmful weights. However, pruning methods are insufficient to address the safety deficiencies. And these approaches focus only on local safety importance, lacking a comprehensive consideration of the internal relationships between safety layers and safety neurons, which limits ability to balance persistent safety and generalizability.
3 Method
@@
Addressing safety concerns in fine-tuned models, we aim to achieve high defense success rates with minimal parameter modifications under a training-free setting, building the capability for continuous safety improvement. To achieve this, we design a multi-scale fine-grained safety activation localization method and perform safety-direction projection on the sparse safety neuron weights, as shown in Fig. 1.
3.1 Safety-Critical Layer Identification
@@
To analyze the varying impacts of different Transformer layers on safety alignment, we first theoretically compared internal states between unaligned and aligned models when responding to benign versus harmful prompts. We used 100 benign prompts from the Alpaca dataset Taori et al. (2023) and 100 harmful prompts from JailbreakBench Mazeika et al. (2024), conducting five repeated experiments with LLaMA 3.1-8B Meta AI (2024). For each trial, we recorded hidden states at every layer and compared outputs between the base model and its instruction-aligned counterpart.
@@
Mean hidden states were collected per layer under four conditions: (1) base model (LLaMA 3.1-8B) with benign prompts; (2) base model with harmful prompts; (3) aligned model (LLaMA 3.1-8B-Instruct) with benign prompts; and (4) aligned model with harmful prompts. The aligned model is defined in Eq. 1, with the base model formulated analogously.
@@
$\begin{array}[]{l}S_{\text{align}}(\text{benign})=\frac{1}{B}\sum_{b=1}^{B}D_{b}(o_{b}^{0},o_{b}^{1},\dots,o_{b}^{k},\dots,o_{b}^{K-1})\\[6.0pt] S_{\text{align}}(\text{harm})=\frac{1}{H}\sum_{h=1}^{H}D_{h}(o_{h}^{0},o_{h}^{1},\dots,o_{h}^{k},\dots,o_{h}^{K-1})\end{array}$ (1)
@@
Here, $o_{b}^{k}$ denotes the hidden state at the $k$-th layer of the LLM in benign dataset. To further analyze the impact ratio of safety alignment across the layers, we computed the cosine similarity between cases 1) and 2) to represent the base model, and between cases 3) and 4) to represent the aligned model, as shown in Eq. 2. Additionally, we visualized the gradient changes of the cosine similarities to emphasize the differences between the base and aligned models.
@@
As shown in Fig. 2a, both models show decreasing cosine similarity across prompts, indicating improved harmful prompt discrimination from shallow to deep layers. Significant differences appear between layers 10–25, where the base model maintains higher similarity, suggesting alignment enhances differentiation capability in mid-deep layers. Gradient analysis (Fig. 2b) shows the sharpest divergence occurs at layers 10–15, where internal representations begin to substantially diverge at about one-third of the model’s depth, causing differing downstream behaviors.
Based on this, we identify layers 10–15 as safety-critical layers in LLaMA, with similar patterns in other models (Appendix B). Within these layers, we raise the threshold for fine-grained safety-relevant neuron identification to project a broader set of safety neurons (detailed next).
Figure 2: Visualization of Safety Activations in Multi-Scale Layers and Neurons. (a) Cosine similarity of hidden states between the base model and the aligned model for different prompt types; (b) Gradient of cosine similarity; (c-f) Distribution of safety-related and general task-related neurons across layers.
3.2 Fine-Grained Safety Neuron Localization
After identifying safety-critical layers, we integrate multi-scale layer information into fine-grained safety neuron localization for enhanced sparsity and precision. Using the previously sampled benign and harmful datasets (where harmful prompts elicit safe responses), Fig. 2c–f visualizes the top 50% highly activated neurons for each type. Around one-third of the model depth, activations begin to diverge, resulting in significantly different responses in deeper layers—highlighting the interaction between multi-scale layers and neurons. Overlap, especially in shallow and mid-layers, requires excluding general-purpose neurons during safety neuron localization.
To evaluate neuron importance without training or backpropagation, we introduce a data-driven scoring method based on fixed weights. The safety-relevant neuron importance score $C_{j}^{harm}$ is computed as Eq. 3, with the general-task neuron score $C_{j}^{benign}$ derived similarly.
@@ 3.2 Fine-Grained Safety Neuron Localization
$\mathcal{Mask}_{l}[j]=\mathbb{I}\left(j\in\text{Top}_{q_{l}}(C^{\text{harm}})\setminus\text{Top}_{p_{l}}(C^{\text{benign}})\right)$ (4)
@@
3.3 Training-Free Sparse Projection
@@
Next, we project the identified sparse safety neurons onto the safety direction. This work employs the LoRA method Hu et al. (2022) for efficient fine-tuning on downstream tasks. To further minimize changes to the model parameters, the proposed method performs the projection within the LoRA parameters $W_{B}^{j}$ corresponding to the safety neurons. The computation is formulated as follow:
@@
3.4 Continual Safety Neuron Projection
@@
Where $\mathbb{I}()$ denotes the indicator function. Sparse heterogeneous projections maintain consistency in safety direction, enhancing the safety of new dimensions while preventing catastrophic forgetting of previously adjusted safety dimensions.
Model Method Edit Param GPT-4o Judger Llama3.1- 405B Judger Keyword ASR AlpacaEval Winrate
@@
Our MSCP 4.67% 1.37 1.36 14% 54.52%
Table 1: Safety and Utility Across Different Defense Methods on Alpaca-Finetuned Models.
Method Judger Score/ASR Param Accuracy
Lora Finetune GPT-4o 3.95 0% 54.20%
@@
4 Experiments
4.1 Experimental Settings
Datasets and Models. To evaluate the effectiveness, we adopt Llama3.1-8B-Instruct Meta AI (2024) and Qwen2.5-7B-Instruct Yang et al. (2024) using the parameter-efficient fine-tuned LoRA approach on two distinct datasets: the instruction-following semantic Alpaca Taori et al. (2023) and the mathematical reasoning GSM8K Cobbe et al. (2021). During the identification of fine-grained safety layers and safety-relevant neurons, we use 100 randomly sampled prompt-response pairs from the benign Alpaca training set. The safety dataset comprises 100 harmful prompts from the “Goal” module of the JailbreakBench Mazeika et al. (2024), along with safe responses generated by GPT-4o OpenAI (2024). For defense evaluation, we assess overall safety performance using 100 randomly sampled general safety prompts from BeaverTails Ji et al. (2023). This dataset also enables multi-dimensional safety evaluation to assess the continual alignment capabilities of MSCP. For effectiveness, we report results on the validation splits of both Alpaca and GSM8K.
Baseline Methods. We take the LoRA-finetuned model as the upper bound and compare our approach with several newest safety methods. Among them, self-reminders Xie et al. (2023) are a prompt-based defense method through system-mode. Goal Priority Zhang et al. (2023) combines discriminator enhancement and fine-tuning by transforming jailbreak prompts into safe refusals through targeted model tuning. SafeLoRA Hsu et al. (2024) and Wanda Wei et al. (2024) are both based on the same post-fine-tuning safety defense paradigm as ours, with SafeLoRA editing safety-relevant layers and Wanda targeting safety-relevant neurons.
Safety CL MSCP CL Judge Method Lora Finetune Universal Safety Animal Abuse Child Abuse Terrorism